OpenStudy (anonymous):

what is the y intercept of 12x+4y=10? A. -3 B. 2.5 C. 4 D. -12

OpenStudy (anonymous):

B is the correct answer, Y intercept means the point where line cuts the y-axis (i-e x=0); Now; 12x+4y = 10 put x=0; 4y = 10 y = 10/4 y=2.5 Hope This helps.

OpenStudy (anonymous):

thank you! @terrenium

OpenStudy (anonymous):

Any time :)